Accelerator Update for December 12, 2007 - December 14, 2007
Wednesday December 12
The day shift began with Operations monitoring the Tevatron (TeV) store 5795 (74.83E30), the Recycler stash (186.1E10), and the Antiproton Source (Pbar) stack (14.9E10), with the Main Injector (MI) and Booster operating normally, but with no beam due to the Linac access to repair the Momentum dump vacuum leak, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, and with NuMI , SciBooNE / MiniBooNE, and the Meson MTest experiments off due to the Linac access.
Operations began preparing Booster for access at 8:24 AM.
Operations began preparing NuMI for access at 8:53 AM.
A Linac expert reported at 11:12 AM that they had installed the vacuum window at the entrance of the Momentum dump. They need to reinstall the BPM and conduct a vacuum leak check before they can begin pumping down. Here is a picture of the vacuum cap.
Booster came out of access at 11:49 AM.
Operations established a MI ramp at 12:11 PM.
Instrumentation experts began work on the MI DCCT at 2:03 PM. The work took about an hour.
A vacuum expert reported at 3:17 PM that the dump side of the Linac vacuum had a bad seal. They will replace the vacuum seal and restart the pump down.
NuMI came out of access at 4:12 PM.
At 8:51 PM, Operations reported that six Booster power supplies in the 400 MeV area lost power. Experts were called in to help investigate.
Operations reported at 9:40 PM that a Linac expert had established beam to the Linac dump.
Operations established beam to Pbar at 9:44 PM.
Operations established beam to NuMI at 9:59 PM.
At 11:04 PM, Operations tried to establish beam to MTest, but a power supply (S:Q202) would not ramp. Operations called in an expert.
Operations reported at midnight that the Recycler had stashed antiprotons once during the evening shift.
Thursday December 13
The midnight shift began with Operations monitoring store 5795 (34.0E30) and the stash (219.8E10), with Pbar stacking (16.5E10), and with NuMI and Sci /MiniBooNE taking beam.
Operations terminated store 5795 at 12:36 AM.
Operations put the TeV into shot setup at 12:48 AM.
At 1 AM, Operations received a smoke alarm for the TeV sector C4 cryo system service building. Operators investigated and found a serious problem with the cold compressor. Operations contacted the Cryo Coordinator. The shaft in this picture should be straight. Shot setup was put on hold for the cold compressor repairs.
Cryo system experts reported at 5:44 AM that they had completed their repairs on the C4 cold compressor. The C4 pump down was completed about an hour later.
Operations put the TeV through a dry squeeze at 6:56 AM.
At 7:08 AM, a TeV sector A17 separator sparked.
Operations put the TeV through a wet squeeze at 8:07 AM.
Operations reported at 8:12 AM that the Recycler had stashed twice during the midnight shift.
At 8:26 AM, the Duty Electrician reported that he needed to access the breaker panel. The following list of devices had to be turned off to power down the panel: LRF2, RF systems 1 & 2, vacuum pump outlet tanks 1 & 2, the 25 HP RF water systems 1 & 2, and a sanitary lift station.
At 9:57 AM, the Duty Electrician reported that this was a very old panel and there were no spare breakers for it, nor were the breakers even made anymore. They will search for other panels like this one around the lab that might be scavenged.
At 11:13 AM, the Duty Electrician reported that he found a similar panel behind the Neutron Therapy Facility (NTF) elevator. He will power the panel down and take the breaker.
At 11:17 AM, Cryo system technicians reported that the belts for the C4 wet engine were loose. They will adjust them.
The Duty Electrician reported at 11:52 AM that he replaced the breaker for the LRF1 water skid. Earlier, a Linac expert had found and began repairing a bad pump for the water skid. The bad pump had caused the breaker trip.
At 11:55 AM, a person reported hearing sparking from vacuum equipment at TeV sector E3. An expert investigated and said that he would replace the cold cathode electronics at E3 and some similar equipment at F4.
Operations put the TeV through a dry squeeze at 1:28 PM.
Operations reported at 1:57 PM that Linac experts had the low energy buncher and LRF2 up and running. They also reported that they had half the water skid for LRF1 on and were working on getting the RF pump hooked up.
Linac experts reported at 2:47 PM that LRF1 was up and running.
Operations put the TeV through a wet squeeze at 3:29 PM.
Operations established beam to MTest at 4:21 PM.
Operations put the TeV through a wet squeeze at 4:29 PM.
Operations put the TeV into shot setup at 5:12 PM. Shot setup holds off sending beam to MTest.
NTF reported at 5:34 PM that they had completed treatment on their last patient for the day.
Operations established store 5800 at 7:49 PM with an initial luminosity of 174.35E30.
Operations established beam to MTest at 8:56 PM.
Operations reported at midnight that the Recycler had stashed twice during the evening shift.
Friday December 14
The midnight shift began with Operations monitoring store 5800 (90.27E30) and a stash (130.3E10), with Pbar stacking (41.0E10), and with MTest, NuMI, and Sci/MiniBooNE taking beam.
Operations disabled beam to MTest at 12:06 AM.
At 12:30 AM, the MI power supply interlock loop tripped off for the first of two times in two hours. It reset.
Operations established beam to MTest at 7:50 AM.
Operations reported at 7:56 AM that the Recycler had stashed five times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 12/14/2007 email Fermilab|
|Security, Privacy, Legal|